OpenStudy (anonymous):

Find the point (x,y) on the graph of y=√x nearest the point (4,0) Can you also please explain "why" you do each step and not just "how" to solve the problem? Thank you!

OpenStudy (anonymous):

The distance formula is: \[d = \sqrt{(x1 - x2)^2 + (y1 - y2)^2}\] But minimizing d^2 will minimize what's inside the square root, so we can rewrite that as \[d^2 = (x1-x2)^2 + (y1 - y2)^2\] plug in your values \[d^2 = (4 - x)^2 + (0 - \sqrt{x})^2\] which gives us \[d^2 = (4-x)^2 + x\] Expand \[d^2 = 16 - 8x + x^2 + x\] which becomes \[d^2 = x^2 - 7x + 16\] Take the derivative of both sides \[2*d*d/dx = 2x - 7\] But since you want the shortest possible distance, d/dx = 0 which indicates a minimum So you have \[0 = 2x-7\] So x = 7/2 Plug that back into your equation for y to get the point.